Find the specific gravity and the density of an object with a dimension of 25cm



x 60cm x 80 cm with a mass of 9000g.

1
Expert's answer
2021-11-19T10:50:57-0500

Let's first find the volume of the object:


"V=0.25\ m\times0.6\ m\times0.8\ m=0.12\ m^3."

Then, we can find the density of the object:


"\rho=\dfrac{m}{V}=\dfrac{9\ kg}{0.12\ m^3}=75\ \dfrac{kg}{m^3}."

Finally, we can find the specific gravity of the object:


"SG=\dfrac{\rho}{\rho_{water}}=\dfrac{75\ \dfrac{kg}{m^3}}{1000\ \dfrac{kg}{m^3}}=0.075"